Ask your own question, for FREE!
Mathematics 14 Online
OpenStudy (anonymous):

A street magician asks a passerby to draw one card from a standard deck, remember it, and then replace it. The magician then shuffles the deck and asks the person to draw a second card. Find the probability that at least one of the two cards was a heart.

OpenStudy (dumbcow):

probability that neither card was a heart = 3/4 *3/4 = 9/16 prob at least 1 was heart = 1 - 9/16

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
3 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!